$ for Savings (from Budget): $                                                                                                              157,131
Match: 3%
Expected Return: 7%
Years: 35
Question Part Two
#1 Based on you’re the money available for savings from your budget, how much money would you have when you turn 65, assuming you earn a 7% annual return on your investments?  
Annual Savings (be sure to include the employer match)
Total Amount at age 65
#2 During retirement, your return drops to 5% annually as you decide to invest more conservatively to protect your principal. How much could you spend each year and not have your money run out before age 95?
Part Three
#1 Assuming you have a retirement savings goal of $3 million at age 65, how much would you need to save monthly to meet that goal if you can earn a 7% annual return on your investments (assume you have nothing saved today and 35 years until retirement)?
Answer #1
#1
a Annual Savings from budget $157,131
b=a*3% Employer Match $4,713.93
c=a+b Annual Savings $161,844.93
Rate Annual return 7%
Nper Number of years of savings 35
Pmt Annual savings $161,844.93
FV Total amount at age 65 $22,372,937.90
#2
Rate Annual return 5%
Nper Number of years in retirement=95-65 30
Pv Amount available at the time of retirement $22,372,937.90
PMT Amount   you can spend each year $1,455,391.72
Part3
#1
Rate Annual monthly return=((1+0.07)^(1/12))-1 0.5654%
Nper Number of Months of savings=35*12 420
Fv Retirement goal $3,000,000
PMT Amount needed to be saved each month $1,752.94